the length of a rectangle is 5 cm more that twice the width. the perimeter of the rectangle is 34 cm. find the demensions of the rectangke.
Answer by checkley77(12844)

L=2W+5
2L+2W=34
2(2W+5)+2W=34
4W+10+2W=34
6W=34-10
6W=24
W=24/6
W=4 ANS. FOR THE WIDTH.
L=2*4+5=8+5=13 ANS. FOR THE LENGTH
PROOF:
2*13+2*4=34
26+8=34
34=34